OpenStudy (anonymous):

y = log x If y = 10, then what is x?

OpenStudy (anonymous):

hello...:)

OpenStudy (anonymous):

hello!

OpenStudy (anonymous):

ok summer, you'll need to solve 10=logx

OpenStudy (apoorvk):

Hello too! :D

OpenStudy (anonymous):

hello to everyone lol and yes i need help!

OpenStudy (anonymous):

can you write that out in it's equaivalent exponential form?

OpenStudy (anonymous):

i have the answer but i need to know how to get it?

OpenStudy (anonymous):

use this...\(\large log_bM=n \Rightarrow b^n=M \)

OpenStudy (anonymous):

answer- x = 1

OpenStudy (anonymous):

ya, all you need to do is write the log equation, \(\large 10=logx\rightarrow 10^{10}=x \)

OpenStudy (anonymous):

wait... how can x=1?

OpenStudy (anonymous):

idk !

OpenStudy (anonymous):

is y=0 or y=10? if y=0 then x=1

OpenStudy (anonymous):

i think it is y=0?

OpenStudy (anonymous):

that's the only way x=1 because then you'll have \(\large 10^0=1 \)

OpenStudy (anonymous):

oh? i did not know that!lol

OpenStudy (anonymous):

could be a misprint on your book? or if the question is written correctly, then x= \(\large10^{10} \)

OpenStudy (anonymous):

ok so what would it be? i am really confused lol sorry i am not that great at math!

OpenStudy (anonymous):

it depends on what question were talking about... if the question is \(\large 0=logx \) then x=1, but if the question is \(\large 10=logx \) then \(\large x=10^{10} \)

OpenStudy (anonymous):

as stated originally, it would be that second one.

OpenStudy (anonymous):

its the first one!

OpenStudy (anonymous):

so y=0 then? and that implies 0=logx ???

OpenStudy (anonymous):

ya?

OpenStudy (anonymous):

ok.... so it is a misprint...y=0 will give you the equation: \[\large 0=logx \rightarrow 10^0=x\rightarrow x=1 \]

OpenStudy (anonymous):

ok that makes more sence! thanks so so so so much!!!

OpenStudy (anonymous):

yw...:)